. If the frequency of two alleles in a gene pool is 90% A and 10%a, what is the
frequency of individuals in the population with the genotype Aa will be____?​
1
Expert's answer
2020-11-10T10:25:50-0500

 If the frequency of two alleles in a gene pool is 90% A and 10%a, what is the

frequency of individuals in the population with the genotype Aa will be 18% (0,18)

p(A)=90%=0,9 in accordance with the Hardy-Weinberg equation:

q(a)=10%=0,1 the frequency of hereozygous pq individuals is 2pq

2pq-? 2pq=2*0,9*0,1=0,18 (18%)